Minute of KAGRA commissioning meeting
- Date: 2018/7/24 11AM~12PM
- Attendance: Miyakawa, Aso, Enomoto, Yokozawa, Kozakai
- eZuce: Miyoki, Suzuki, Nakano, Shoda, Ushiba, Yamamoto, Kimura, Uchiyama
- 40W laser started working on last Friday.
- Limite 1W will be used to alignment or so.
- Remote desktop will be installed on Windows PC for 40W laser control.
- Monitor of laser power should be displayed on door or something.
- MZ of single arm mode matching is much better now. No Lager-Gaussian is seen.
- MZ contrast is 99.4% and 96.9%. Second one has 3m arm length difference.
- Lock wil be tried soon. Common mode servo will be used to control MZ.
- RAM(residual amplitude monitor) monitor will be installed.
- ISS TF model and measurement were done. Model still does not have phase delay at high frequency.
- RIN is limited by PD dark noise up to 1kHz. Shotnoise is almost comparable.
- UGF is currently 10kHz. Above 1kHz, suppression is not enough. It will be limited to 50kHz max in future by a commercial AOM driver of speed of variable attenuator..
Check list for IXC
- Meeting will be assigned on one of 8/15,16,17.
- WAB, NAB are the additional items.
- Mirror position should be in the item list.
- WAB PD will be installed at the sometime.
- The number of opening flange should be minimized. It must be listed up.
- Oplev of ITMX will be installed.
- Safety for high power laser must be considered.
- Oplev power goes down with switch of cryostat.
- Also ETMX QPD shows 1/10 or lower power in 10 hour without relationship of switch of cryostat. One of the viewport looks dewed.
- One of the motor seemed to be stuck.
- F2 of ETMX fell bistable point, and it is stuck.
- BS is almost done.
- PEM meeting will be held on 8/1 6pm.
- Next meeting: 8/7.